When I am as old as my father is now, I shall be five times as old as my son is now. By then my son will be eight years older than I am now. The combinded ages of my father and myself are 100 years. How old is my son?
Father=x Myself=y Son=s
x=5s x+y=100 5s=y+8 y/5=s
5(y/5)+y=100

I just don't know how to solve this. Please help. Thank you




Answer by rapaljer(4671)

Your equations:
x=5s x+y=100 5s=y+8 y/5=s
5(y/5)+y=100

I think your equations should be:
x=5s x+y=100 s+(x-y)=y+8 x/5=s
y=100-x s+ x - 2y=8
y=100 - 5s

s+x-2y = 8
s+ 5s -2(100-5s) = 8
6s -200 +10s = 8
16s -200=8
16s = 208
s= 13 son's age now
x= 5(13) = 65 father's age now
y=100-x = 100-65 = 35 my age now

Check:
"When I am 65, I shall be five times as old as my son is now." ---- True!!

"By then (I went from 35 to 65, so 30 years has gone by!), my son will be 8 years older than I am now." Son will be 13 + 30 = 43, which is 8 years older than I am now! ---- True!!

"The combined ages of my father and myself are 100 years." ---- True!!

R^2 at SCC